OpenStudy (anonymous):

Solve 8(3a+6)=9(2a-4) in steps?

OpenStudy (anonymous):

<=>24a + 48 = 18a -36 <=> 24a -18a = -48 - 36 <=> 6a = -84 <=> a = -84/6 = -14